OpenStudy (anonymous):

how do i solve this im so confused @.@ Solve the following equation: x^2 - 14x + 49 = 0 {7} {-7} {-7, 7} { 7,-7

OpenStudy (whpalmer4):

You have a number of options. Because this is a multiple-guess question, you could simply try out the answers and see which one works. For example, (7)^2 - 14(7) + 49 = 49 + 49 - 98 = 0, so 7 is a solution. How about -7? (-7)^2 - 14(-7) + 49 = 49 + 98 + 49 = 196, so -7 is not a solution. Now, that works, but only if you have a short menu of answers from which to choose. Better to know how to solve it! \[x^2-14+49 = 0\]Let's try factoring. What are two numbers that when multiplied give us 49, and when added, give us -14?

OpenStudy (anonymous):

ill be honest just use mathway.com it will help you out alot

OpenStudy (whpalmer4):

@chancehenninger Not when she's taking a test, it won't.

OpenStudy (anonymous):

-7?

OpenStudy (whpalmer4):

@EmmaH are you there?

OpenStudy (anonymous):

my teachers always said its not cheating its using your resources

OpenStudy (whpalmer4):

Right! So we can write our equation as \[(x-7)(x-7) = 0\]because \[(x-7)(x-7) = x^2-7x-7x+49 = x^2-14x+49\] Now can you find the values of \(x\) that make \[(x-7)(x-7) = 0\] true?

OpenStudy (anonymous):

your just confusing her even more

OpenStudy (whpalmer4):

@chancehenninger Who said anything about cheating. I'm making sure Emma can solve the problem herself, so that she is her own best resource.

OpenStudy (whpalmer4):

If you have a product of two numbers = 0, one or both of those numbers must be 0. For \[(x-7)(x-7) = 0 \rightarrow (x-7) = 0\]\[x-7 = 0\]\[x=7\] So both solutions to the quadratic equation are x = 7. If you get an equation that you can't or don't want to factor, you can use the quadratic formula to find the solutions of \(ax^2+bx+c=0\) \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]so long as \(a\ne0\) In this case, we have \(a=1,b=-14,c=49\): \[x=\frac{-(-14)\pm\sqrt{(-14)^2-4(1)(49)}}{2(1)} = \frac{14\pm\sqrt{196-196}}{2} = 7\pm 0 = 7 \]

OpenStudy (anonymous):

ok im sorry but (x-7) = 0\]\[x-7 = 0\]\[x=7\] is just a bunch o =f symnbols to me it makes no sense

OpenStudy (whpalmer4):

Well, if I tell you that I have two numbers, and I multiply them together, and the answer is 0, what can you tell me about those numbers?

OpenStudy (anonymous):

all im seing is atuff like this = 0\]

OpenStudy (whpalmer4):

Oh, it isn't displaying your page properly :-( Try refreshing the page. Here, I'll attach pictures of the posts, properly displayed:

OpenStudy (whpalmer4):

All right, then factoring is probably the only approach for you now. Do you see how it produced the answer?

OpenStudy (anonymous):

can you show me how you got it from the beginnign steps im understanding the last steps but not the furst couple

OpenStudy (whpalmer4):

The factoring part? I'll do another one. This one will have two different solutions instead of one. x^2 + 3x + 2 = 0 (I won't do the formatting this time) To factor that, we need to find 2 numbers that when we multiply them give us 2, but when we add them give us 3. Here's why: (x+a)(x+b) = x^2 + ax + bx + ab = x^2 + (a+b)x + ab Now if we look at our equation we are trying to factor, we can see that those two look identical so long as ab = 2 and (a+b) = 3, right?

OpenStudy (whpalmer4):

I'll take that as a yes :-) 1*2 = 2 1+2 = 3 So we can write our factored equation like this: x^2 + 3x + 2 = (x+1)(x+2) = 0 Now, for that to be true, there are two possible values of x: (x+1) = 0 (x+2) = 0 and those give us x = -1, x = -2 as our solutions. Let's check them: (-1)^2 + 3(-1) + 2 = 1 -3 + 2 = 0 (-2)^2 + 3(-2) + 2 = 4 - 6 + 2 = 0